# Learning from Heterogeneity: A Dynamic Learning Framework for Hypergraphs

### v1

@@ V-F Impact of Hyperparameter $\alpha$
In this paper, we propose a heterogenous hypergraph learning framework for representation learning. This framework first generates the high-quality initial node embedding using the designated pairwise fusion function, aiming to exploit the pairwise graph information at most. Afterwards, the multi-type hyperedges are constructed dynamically, forming the hypergraph together. The embedding is then updated iteratively through type-specific attention, aiming to encode the heterogeneous attributes into the embedding space. We conduct comprehensive experiments on three widely-used public datasets and comparison with eleven baseline methods to demonstrate the effectiveness of our proposed framework. The results and analysis reveal that the proposed framework can achieve new state-of-the-art performance on both node classification and link prediction tasks.

### v2

@@ V-G Visualization
In this paper, we propose a heterogenous hypergraph learning framework for representation learning. This framework first generates the high-quality initial node embedding using the designated pairwise fusion function, aiming to exploit the pairwise graph information at most. Afterwards, the multi-type hyperedges are constructed dynamically, forming the hypergraph together. The embedding is then updated iteratively through type-specific attention, aiming to encode the heterogeneous attributes into the embedding space. We conduct comprehensive experiments on three widely-used public datasets and comparison with twelve baseline methods to demonstrate the effectiveness of our proposed framework. The results and analysis reveal that the proposed framework can achieve new state-of-the-art performance on both node classification and link prediction tasks.